OpenStudy (anonymous):

1. determine the distance between twogiven lines > L1 2x - y = 5 > L2 4x - 2y = 7 2. determine the equation of the line passing through the given point ( -1,-1) and perpendicular to x -y + 3 = 0 3. determine the equation of the line passing through the given point (-2 , 5) and parallel to x + 3y - 5 = 0 FORMULAS: absolute value of AX BY +C all over 2a (short method) square root of (X1+X2)(Y1+Y2) (long method)

OpenStudy (dumbcow):

For L1: the y-intercept is -5 x=0 -> -y =5 -> y=-5 For L2: the y-intercept is -3.5 x=0 -> -2y = 7 -> y = -3.5 distance = |-5 -(-3.5)| = 1.5 is that what you are looking for?

OpenStudy (anonymous):

2) x -y + 3 = 0 x+3 = y perpendicular slope = -1 y- (-1)= -1(x-(-1) ) y+1 = -x -1 y = -x-2

OpenStudy (anonymous):

3) x + 3y - 5 = 0 rewrite slope intercept y = (-x +5)/3\[y=-\frac{1}{3}x+\frac{5}{3}\] slope = -1/2 parallel than slope is the same -1/2 \[y-5= -\frac{1}{3}(x-(-2))\]\[y=-\frac{1}{2}x+\frac{13}{3}\]

OpenStudy (anonymous):

number 1 ) it been so long I don't renember

OpenStudy (anonymous):

tnx2.. ;D

OpenStudy (anonymous):

np, number 1) it not hard , I don't have time go back to look the lesion

OpenStudy (anonymous):

But you got answer number 1

OpenStudy (dumbcow):

i might be wrong, i gave the vertical distance between the lines or the vertical shift but Not the shortest distance between the 2 lines. To do that you need to find a perpendicular line that goes through both lines L1,L2 then use distance formula to find distance between the intersection points

OpenStudy (anonymous):

number I no idea , don't renember

OpenStudy (anonymous):

number 1) no idea